Two blocks of mass 300 g and 200 g respectively are moving toward each other along a horizontal frictionless surface with velocities of 50 cm/s and 100 cm/s, respectively. a. If the blocks collide and stick together, find their final velocity. (5 pts.) b. Find the final velocity of each block if the collision is completely elastic (e =1).


1
Expert's answer
2021-05-26T21:35:25-0400

Gives

"m_1=300g=0.3kg"

"m_2=200g=0.2kg"

"u_1=50cm\/sec=0.50cm\/sec"

"u_2=100cm\/sec=1m\/sec"

"v=?"

Momentum conservation

"m_1u_1+m_2u_2=(m_1+m_2)v"

"v=\frac{m_1u_1+m_2u_2}{m_2+m_2}"


"v=\frac{0.3\times0.50+0.2\times1}{0.3+0.2}=0.7m/sec"

(b)

Momentum conservation in elastic collision

"m_1u_1+m_2u_2=(m_1v_1+m_2v_2)"


"0.3\times0.5-0.2\times1=0.3v_1+0.2v_2"

"0.3v_1+0.2v_2=0.15-0.2"

"0.3v_1+0.2v_2=-0.05\rightarrow(1)"

elastic collision

e=1

"v_2-v_1=u_1-u_2"

"v_2-v_1=0.5+1=1.5\rightarrow(2)"

eqution (1)and(2)

"v_2=80cm\/sec"

"v_1=-70cm\/sec"